OpenStudy (chinaramirez173):

Given the function f(x) = 0.5|x – 4| – 3, for what values of x is f(x) = 7? A: x = –24, x = 16 B: x = –16, x = 24 C: x = –1, x = 9 D: x = 1, x = –9

OpenStudy (welshfella):

you could either plug in the given values to see which ones fit or solve the equation 7 = 0.5|x – 4| – 3 for x

OpenStudy (sshayer):

0.5|x-4|-3=7 0.5|x-4|=7+3=10\[\left| x-4 \right|=\frac{ 10 }{ 0.5 }=10 \times \frac{ 10 }{ 5 }=20\] \[x-4=\pm20\] find x

OpenStudy (welshfella):

0.5|x – 4| = 10 |x – 4| = 20 can you continue from here?

OpenStudy (chinaramirez173):

i havent learned this yet so i dont really know how to solve it

OpenStudy (welshfella):

the vertical lines around the x - 4 mean the Absolute value of x - 4 which is positive even if x - 4 = negative so we split this up as sshayer did x - 4 = 20 or x - 4 = -20 can you solve these 2 equations now?

OpenStudy (welshfella):

all you have to do is isolate x by adding 4 to both sides of the equations

OpenStudy (welshfella):

x - 4 + 4 = 20 + 4 x = 24 thats one answer can you do the other one?

OpenStudy (welshfella):

first line is x - 4 + 4 = -20 + 4

OpenStudy (chinaramirez173):

i got 6 for x-4=20

OpenStudy (welshfella):

no - i did that one for you already

OpenStudy (welshfella):

look at my 5th post

jhonyy9 (jhonyy9):

so for you get a value of 20 resulted from this modulus so you need giving to x or 24 because 24-4=20 and multiplied 20 by 0,5 = 10 and 10 -3 = 7 what you need get it for f(x) and in the second case than you give to x the value of -16 so will get inside this modulus -16-4 = -20 what is resulted +20 bc. in modulus so 0,5 multiplied by 20 710 and 10-3=7 so f(x) = 7 hope this is now more understandably good luck

jhonyy9 (jhonyy9):

so choice B. is right sure